(1 point) Solve y + 2y + 2y = 4te* cos(t). 1) Solve the homogeneous part: y + 2y + 2y = 0 for Yh, using a real basis. Not
Answer #1

4 + 24 + 2y=uté & cost The AE is 22 +227950 ar=-1 Ii ① Y = cilet sin(t) + G (et collt)) Answer ♡ natural for grencies are iti® W+3942= A exp ((+1+i) )(-ait+6+4+41)t ta) Answer Hence A=tli Bari Brelyy) = test cost) - tetsinlt)+ té cosct) ttet sin(t)
Answer #2

you + xy + ay = até + corte) -0 () Consider the homogeneous equation y + ey + 2y =0 Characteristic (auxiliary) equation is:= (, costo + C, diver) y G c* coste +ge tainty (2) Comides et tilt , tr ( cao (1) tủ 4 (1) E et conlt) ti ét ain(t) :. et co% = 16€**}* ) (A++8) = (A4° +84)861+65+ Yo = (24+ +6) EL**)* + (at +62) 2*) (1-1) = (24+8) <++i) +(-At? - 8t + Ait? + Bit)E= elitist (-dac)t? + (+-4A – 366)+ +(29 -26+81) Coneider Yo + 2y/ + og = CM2)t | (+264) £® + (4 10 –44-20i)t + (24-28+381)]:. + 24 + 37p = 4+el+) → Rel++)! ( fi ++ (av 341) - 4tectie 2A(46+ + (444)) - 4+ => 41A = 4 and 1(2+ 381)=0 - Ai =L =) Art.. A = B = 1 % = + (&?*0)%) (+it+2) + + -L-ittt) ** (corte) t i din te) =€* (-it +1) co lt) + il-ittt) e *rint) =+* [-it.